# Slender Trees and the Approximation Property

### v1

@@ 7. Making $\operatorname{ISP}$ indestructible under directed closed forcing
Let $\kappa$ be a cardinal and $l\colon\kappa\to V_{\kappa}$. $l$ is a Laver function if for any $A$ and any $\lambda\geq|\operatorname{tcl}(A)|$ there is an embedding $j\colon V\to M$ with critical point $\kappa$ such that $j(\kappa)>\lambda$, $j(f)(\kappa)=A$ and $M^{\lambda}\subseteq M$.
@@
Theorem 7.2.
Let $\kappa$ be a supercompact cardinal. Then there exists a Laver function $l\colon\kappa\to V_{\kappa}$
@@
Let $\kappa$ be a supercompact cardinal. Then the following holds: For any $\lambda\geq\kappa$, any $(\kappa,\lambda)$-list $f$, any $A\in H(\lambda^{+})$ and any large enough $\Theta$ there exists a $\lambda$-ineffability witness $N$ for $\kappa$ with respect to $f$ such that, letting $\pi$ be the Mostowski-Collapse of $N$, $l(\pi(\kappa))=\pi(A)$.
@@
Let $f$ be a $(\kappa,\lambda)$-list. Let $A\in H(\lambda^{+})$, $\Theta$ be large and $x\in[H(\Theta)]^{<\kappa}$. Let $\delta:=|H(\lambda^{+})|$ and $j\colon V\to M$ an elementary embedding such that $j(\kappa)>\delta$, ${}^{\delta}M\subseteq M$ and $j(l)(\kappa)=A$. By the closure properties $j[\lambda]\in M$, so $b:=j^{-1}[j(f)(j[\lambda])]$ is defined. Inside $V$, let $\Theta$ be large. Then the set $C^{\prime}$ of all $N\prec H(\Theta)$ of size ${<}\,\kappa$ containing $b,f,\kappa,\lambda,A$ is a club. Hence the set
$C:=\{N\cap H(\lambda^{+})\;|\;N\in C^{\prime}\}$
contains a club, so $j[H(\lambda^{+})]\in j(C)$. Thus there exists $N\prec H(j(\Theta))^{M}$ containing $j(b),j(f),j(\kappa),j(\lambda),j(A)$ with $N\cap j(H(\lambda^{+}))=j[H(\lambda^{+})]$. In particular, $N\cap j(\lambda)=j[\lambda]$ and $N\cap j(A)=j[A]$, so
@@
Note first that the statement makes sense because if $x\in H(\lambda^{+})$, $j(x)\in j[H(\lambda^{+})]\subseteq N$. We prove it by $\in$-induction. Assume $y\in H(\lambda^{+})$ and $\pi(j(x))=x$ for every $x\in y$. By the definition
@@
Assume $a\in j(y)\cap N$. Because $j(y)\in j(H(\lambda^{+}))$, $a\in N\cap j(H(\lambda^{+}))=j[H(\lambda^{+})]$, so there is $x\in H(\lambda^{+})$ with $j(x)=a$. By assumption $x\in y$ so $\pi(j(x))=x\in\pi(j(y))$. Assume $x\in y$. Then $j(x)\in j(y)$ and $j(x)\in j[H(\lambda^{+})]$, so $j(x)\in N$. Thus $\pi(j(x))=x\in\pi(j(y))$. ∎
Thus $j(l)(\pi(j(\kappa)))=j(l)(\kappa)=A=\pi(j(A))$. In summary, in $M$, there exists a $j(\lambda)$-ineffability witness (namely $N$) for $j(\kappa)$ with respect to $j(f)$ such that $j(l)(\pi(j(\kappa)))=\pi(j(A))$. By elementarity there is in $V$ a $\lambda$-ineffability witness for $\kappa$ with respect to $f$ such that $l(\pi(\kappa))=\pi(A)$. ∎
The previous lemma allows us to use small embeddings to prove our last result. For technical reasons, let $\operatorname{Add}(\tau,\beta)$ consist of functions $f$ on $\beta$ of size ${<}\,\tau$ such that for all $\gamma\in\operatorname{dom}(f)$, $\gamma$ is a successor ordinal and $f(\gamma)\in\operatorname{Add}(\tau)$.
@@
    $q$ is a partial function on $\beta$ of size ${<}\,\mu$ such that if $\gamma\in\operatorname{dom}(q)$, $\gamma=\delta+2$ for an inaccessible cardinal $\delta$ and $q(\gamma)$ is an $\operatorname{Add}(\tau,\gamma)$-name for a condition in $\dot{\operatorname{Coll}}(\check{\mu},\check{\delta})$
@@
    $r$ is a partial function on $\beta$ of size ${<}\,\mu$ such that if $\gamma\in\operatorname{dom}(r)$, $\gamma$ is inaccessible, $l(\gamma)$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma)$-name for a ${<}\,\gamma$-directed closed partial order and $r(\gamma)$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma)$-name for an element of $l(\gamma)$.
@@
    $\mathbb{M}_{3}^{l}(\tau,\mu,\nu)$ is ${<}\,\tau$-directed closed,
@@
    The term ordering on $\mathbb{M}_{3}^{l}(\tau,\mu,\nu)$ is ${<}\,\mu$-closed,
@@
The proofs for (1), (2) and (3) are standard. Regarding (4), let $(p,q_{\alpha},r_{\alpha})_{\alpha<\delta}$ be a descending sequence (with $\delta<\mu$). We first define $q$ as in Lemma 5.2, so that $(p,q)$ is a lower bound of $(p,q_{\alpha})_{\alpha<\delta}$. Now let $y:=\bigcup_{\alpha<\delta}\operatorname{dom}(r_{\alpha})$, which has size ${<}\,\mu$. We define a function $r$ on $y$ by induction on $\beta<\nu$ such that $(p\upharpoonright\beta,q\upharpoonright\beta,r\upharpoonright\beta)$ is a lower bound of $(p\upharpoonright\beta,q\upharpoonright\beta,r_{\alpha}\upharpoonright\beta)_{\alpha<\delta}$. Assume $r\upharpoonright\beta$ has been defined and $\beta\in y$. Let $\alpha_{0}$ be such that $\beta\in\operatorname{dom}(r_{\alpha_{0}})$. For $\alpha_{0}\leq\alpha<\alpha^{\prime}$, $(p,q_{\alpha^{\prime}},r_{\alpha^{\prime}})\leq(p,q_{\alpha},r_{\alpha})$ and thus
@@
so $(p\upharpoonright\beta,q\upharpoonright\beta,r\upharpoonright\beta)\Vdash r_{\alpha^{\prime}}(\beta)\leq r_{\alpha}(\beta)$. Thus $(p\upharpoonright\beta,q\upharpoonright\beta,r\upharpoonright\beta)$ forces $(r_{\alpha}(\beta))_{\alpha_{0}\leq\alpha<\delta}$ to be a descending sequence in some ${<}\,\beta$-directed closed partial order (where $\beta\geq\mu$) and we can fix a lower bound $r(\beta)$. Then $(p\upharpoonright\beta+1,q\upharpoonright\beta+1,r\upharpoonright\beta+1)$ is a lower bound of $(p\upharpoonright\beta+1,q_{\alpha}\upharpoonright\beta+1,r_{\alpha}\upharpoonright\beta+1)_{\alpha<\delta}$. Now we show (5). Regarding the projection property, let $(p^{\prime},q^{\prime},r^{\prime})\leq(p,q,r)$. Find $q^{\prime\prime}$ as in Lemma 5.2 such that $(p,q^{\prime\prime})\leq(p,q)$ and $(p^{\prime},q^{\prime\prime})\leq(p^{\prime},q^{\prime})\leq(p^{\prime},q^{\prime\prime})$. We define a function $r^{\prime\prime}$ with domain $\operatorname{dom}(r^{\prime})$ by induction such that for every $\beta$,
@@
The refinement property is clear. Regarding the mixing property, let $(p,q_{0},r_{0}),(p,q_{1},r_{1})\leq(p,q,r)$. Choose $p_{0},p_{1}\leq p$ such that $p_{0}(1)$ and $p_{1}(1)$ are incompatible and find $q^{\prime}$ such that $(p,q^{\prime})\leq(p,q)$ and $(p_{i},q^{\prime})\leq(p,q_{i})$. It follows that for any inaccessible cardinal $\gamma$, $(p_{0}\upharpoonright\gamma,q^{\prime}\upharpoonright\gamma)$ and $(p_{1}\upharpoonright\gamma,q^{\prime}\upharpoonright\gamma)$ are incompatible. Define a function $r^{\prime}$ on $\operatorname{dom}(r_{0})\cup\operatorname{dom}(r_{1})$ by induction on $\beta$ such that $(p\upharpoonright\beta,q^{\prime}\upharpoonright\beta,r^{\prime}\upharpoonright\beta)\leq(p,q\upharpoonright\beta,r\upharpoonright\beta)$ and $(p_{i}\upharpoonright\beta,q^{\prime}\upharpoonright\beta,r^{\prime}\upharpoonright\beta)\leq(p\upharpoonright\beta,q_{i}\upharpoonright\beta,r_{i}\upharpoonright\beta)$. Assume $\beta\in\operatorname{dom}(r_{0})\cup\operatorname{dom}(r_{1})$ (if $\beta$ is only in one domain, set $r_{i}(\beta)=\emptyset$ for the other $i$). By the inductive hypothesis, $(p_{i}\upharpoonright\beta,q^{\prime}\upharpoonright\beta,r^{\prime}\upharpoonright\beta)$ are incompatible for $i\in 2$, so we can choose a name $r^{\prime}(\beta)$ that is forced by $(p_{0}\upharpoonright\beta,q^{\prime}\upharpoonright\beta,r^{\prime}\upharpoonright\beta)$ to be equal to $r_{0}(\beta)$ and by conditions incompatible with $(p_{0}\upharpoonright\beta,q^{\prime}\upharpoonright\beta,r^{\prime}\upharpoonright\beta)$ to be equal to $r_{1}(\beta)$. Then $(p\upharpoonright\beta+1,q^{\prime}\upharpoonright\beta+1,r^{\prime}\upharpoonright\beta+1)\leq(p\upharpoonright\beta+1,q\upharpoonright\beta+1,r\upharpoonright\beta+1)$ and $(p_{i}\upharpoonright\beta+1,q^{\prime}\upharpoonright\beta+1,r^{\prime}\upharpoonright\beta+1)\leq(p\upharpoonright\beta+1,q_{i}\upharpoonright\beta+1,r_{i}\upharpoonright\beta+1)$. ∎
@@
Let $\nu<\kappa$ be an inaccessible cardinal such that $l\upharpoonright\nu\colon\nu\to V_{\nu}$. Let $\xi:=\nu+1$ and $G$ be $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-generic. In $V[G]$, define the partial order $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi,\beta)$ by induction on $\beta\geq\xi$, letting $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi,\xi):=\{\emptyset\}$ (this defines a name $\mathbb{M}_{3}^{l}(\Gamma,\tau,\mu,\kappa\smallsetminus\xi,\beta)$) Assume that for all $\gamma<\beta$, $\mathbb{M}_{3}^{l}(\Gamma,\tau,\mu,\kappa\smallsetminus\xi,\gamma)$ has been defined and there is a dense embedding from $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma)$ into $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)*\mathbb{M}_{3}^{l}(\Gamma,\tau,\mu,\kappa\smallsetminus\xi,\gamma)$. Then we let $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi,\beta)$ consist of tripels $(p,q,r)$ such that
@@
    $q$ is a ${<}\,\mu$-sized partial function on $\beta\smallsetminus\xi$ such that for all $\gamma\in\operatorname{dom}(q)$, $\gamma$ is the double successor of an inaccessible cardinal $\delta$ and $q(\gamma)$ is an $\operatorname{Add}(\tau,\gamma\smallsetminus\delta)$-name for an element in $\dot{\operatorname{Coll}}(\check{\mu},\check{\delta})$.
@@
    $r$ is a ${<}\,\mu$-sized partial function on $\beta\smallsetminus\xi$ such that for all $\gamma\in\operatorname{dom}(r)$, $\gamma$ is inaccessible, $l(\gamma)$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma)$-name for a $<\gamma$-directed closed partial order and $r(\gamma)$ is an $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi,\gamma)$-name for an element in $l(\gamma)^{G}$ (reimagining $l(\gamma)$ as an $\mathbb{M}(\xi)*\mathbb{M}_{3}(\Gamma,\kappa\smallsetminus\xi,\gamma)$-name).
@@
We do the proof by induction on $\beta\in[\xi,\kappa]$ with the base case being clear. Assume that such an embedding exists for all $\gamma<\beta$. So in particular, the $\mathbb{M}_{3}^{l}(\Gamma,\tau,\mu,\kappa\smallsetminus\xi,\beta)$ is defined. We define the following embedding:
@@
where $\overline{q}$ and $\overline{r}$ are defined as follows: $\overline{q}$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-name for a function such that $\Vdash\operatorname{dom}(\overline{q})=\check{\operatorname{dom}(q)\smallsetminus\xi}$ and for any $\gamma\in\operatorname{dom}(q)\smallsetminus\xi$, $\overline{q}(\check{\gamma})$ is an $\operatorname{Add}(\tau,\xi)$-name (and thus an $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-name) for an $\operatorname{Add}(\tau,\gamma\smallsetminus\xi)$-name corresponding to the $\operatorname{Add}(\tau,\gamma)$-name $q(\xi)$. $\overline{r}$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-name for a function such that $\Vdash\operatorname{dom}(\overline{r})=\check{\operatorname{dom}(r)\smallsetminus\xi}$ and for $\gamma\in\operatorname{dom}(r)\smallsetminus\xi$, $\overline{r}(\check{\gamma})$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-name for an $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma\smallsetminus\xi)$-name corresponding to the $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma)$-name $r(\gamma)$ (using the inductive hypothesis). It follows as in Lemma 5.6 that $\iota$ is a dense embedding. ∎
@@
    $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$ is ${<}\,\tau$-directed closed,
@@
    The term ordering on $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$ is ${<}\,\mu$-closed,
@@
Let $\xi:\nu+1$ for an inaccessible cardinal $\nu$. Let $G$ be $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-generic. In $V[G]$, let $\dot{\mathbb{L}}$ be an $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$-name for a ${<}\,\kappa$-directed closed partial order. Then $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)*\dot{\mathbb{L}}$ has the ${<}\,\tau^{+}$-approximation property.

### v2

@@ 7. Indestructibility of $\operatorname{ISP}$
Let $\kappa$ be a cardinal and $\colon\kappa\to V_{\kappa}. ll$ is a Laver function if for any $A$ and any $\lambda\geq|\operatorname{tcl}(A)|$ there is an embedding $j\colon V\to M$ with critical point $\kappa$ such that $j(\kappa)>\lambda$, $j(f)(\kappa)=A$ and $M^{\lambda}\subseteq M$.
@@
Theorem 7.2 ([15]).
Let $\kappa$ be a supercompact cardinal. Then there exists a Laver function $l\colon\kappa\to V_{\kappa}$.
@@
Let $\kappa$ be a supercompact cardinal and let $l\colon\kappa\to V_{\kappa}$ be a Laver function. Then the following holds: For any $\lambda\geq\kappa$, any $(\kappa,\lambda)$-list $f$, any $A\in H(\lambda^{+})$ and any large enough $\Theta$ there exists a $\lambda$-ineffability witness $N\prec H(\Theta)$ for $\kappa$ with respect to $f$ such that, letting $\pi$ be the Mostowski-Collapse of $N$, $l(\pi(\kappa))=\pi(A)$.
@@
Let $f$ be a $(\kappa,\lambda)$-list. Let $A\in H(\lambda^{+})$, $\Theta$ be large and $x\in[H(\Theta)]^{<\kappa}$. Let $\delta:=|H(\Theta)|$ and $j\colon V\to M$ an elementary embedding such that $j(\kappa)>\delta$, $^{\delta}M\subseteq M$ and $j(l)(\kappa)=A$. By the closure properties $j[\lambda]\in M$, so $b:=j^{-1}[j(f)(j[\lambda])]$ is defined. The set $C$ of all $N\prec H(\Theta)$ of size $\kappa$ containing $b,f,\kappa,\lambda,A$ is a club. Hence $N:=j[H(\Theta)]$ is in $j(C)$. We have
@@
Note first that the statement makes sense because if $x\in H(\Theta)$, $j(x)\in N$.
We prove it by $\in$-induction. Assume $y\in H(\Theta)$ and $\pi(j(x))=x$ for every $x\in y$. By the definition
@@
Assume $a\in j(y)\cap N$. Because $j(y)\in j(H(\Theta))$, $a\in N\cap j(H(\Theta))=j[H(\Theta)]$, so there is $x\in H(\Theta)$ with $j(x)=a$. By assumption $x\in y$ so $\pi(j(x))=x\in\pi(j(y))$.
Assume $x\in y$. Then $j(x)\in j(y)$ and $j(x)\in N$. Thus $\pi(j(x))=x\in\pi(j(y))$. ∎
Thus $j(l)(\pi(j(\kappa)))=j(l)(\kappa)=A=\pi(j(A))$. In summary, in $M$, there exists a $j(\lambda)$-ineffability witness (namely $N$) for $j(\kappa)$ with respect to $j(f)$ such that $j(l)(\pi(j(\kappa)))=\pi(j(A))$.
By elementarity there is in $V$ a $\lambda$-ineffability witness for $\kappa$ with respect to $f$ such that $l(\pi(\kappa))=\pi(A)$. ∎
The previous lemma allows us to use small embeddings to prove our last result.
For technical reasons, let $\operatorname{Add}(\tau,\beta)$ consist of functions $f$ on $\beta$ of size $\tau$ such that for all $\gamma\in\operatorname{dom}(f)$, $\gamma$ is a successor ordinal and $f(\gamma)\in\operatorname{Add}(\tau)$.
@@
    $q$ is a partial function on $\beta$ of size $\mu$ such that if $\gamma\in\operatorname{dom}(q)$, $\gamma=\delta+2$ for an inaccessible cardinal $\delta$ and $q(\gamma)$ is an $\operatorname{Add}(\tau,\gamma)$-name for a condition in $\dot{\operatorname{Coll}}(\check{\mu},\check{\delta})$
@@
    $r$ is a partial function on $\beta$ of size $\mu$ such that if $\gamma\in\operatorname{dom}(r)$, $\gamma$ is inaccessible, $l(\gamma)$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma)$-name for a $\gamma-$directed closed partial order and $r(\gamma)$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma)$-name for an element of $l(\gamma)$.
@@
    $\mathbb{M}_{3}^{l}(\tau,\mu,\nu)$ is $\tau-$directed closed,
@@
    The term ordering on $\mathbb{M}_{3}^{l}(\tau,\mu,\nu)$ is $\mu-$closed,
@@
The proofs for (1), (2) and (3) are standard.
Regarding (4), let $(p,q_{\alpha},r_{\alpha})_{\alpha<\delta}$ be a descending sequence (with $\delta<\mu$). We first define $q$ as in Lemma 5.2, so that $(p,q)$ is a lower bound of $(p,q_{\alpha})_{\alpha<\delta}$. Now let $y:=\bigcup_{\alpha<\delta}\operatorname{dom}(r_{\alpha})$, which has size $\mu$. We define a function $r$ on $y$ by induction on $\beta<\nu$ such that $(p\upharpoonright\beta,q\upharpoonright\beta,r\upharpoonright\beta)$ is a lower bound of $(p\upharpoonright\beta,q\upharpoonright\beta,r_{\alpha}\upharpoonright\beta)_{\alpha<\delta}$. Assume $r\upharpoonright\beta$ has been defined and $\beta\in y$. Let $\alpha_{0}$ be such that $\beta\in\operatorname{dom}(r_{\alpha_{0}})$. For $\alpha_{0}\leq\alpha<\alpha^{\prime}<\delta$, $(p,q_{\alpha^{\prime}},r_{\alpha^{\prime}})\leq(p,q_{\alpha},r_{\alpha})$ and thus
@@
so $(p\upharpoonright\beta,q\upharpoonright\beta,r\upharpoonright\beta)\Vdash r_{\alpha^{\prime}}(\beta)\leq r_{\alpha}(\beta)$. Thus $(p\upharpoonright\beta,q\upharpoonright\beta,r\upharpoonright\beta)$ forces $(r_{\alpha}(\beta))_{\alpha_{0}\leq\alpha<\delta}$ to be a descending sequence in some $\beta-$directed closed partial order (where $\beta\geq\mu$) and we can fix a lower bound $r(\beta)$. Then $(p\upharpoonright\beta+1,q\upharpoonright\beta+1,r\upharpoonright\beta+1)$ is a lower bound of $(p\upharpoonright\beta+1,q_{\alpha}\upharpoonright\beta+1,r_{\alpha}\upharpoonright\beta+1)_{\alpha<\delta}$.
Now we show (5). Regarding the projection property, let $(p^{\prime},q^{\prime},r^{\prime})\leq(p,q,r)$. Find $q^{\prime\prime}$ as in Lemma 5.2 such that $(p,q^{\prime\prime})\leq(p,q)$ and $(p^{\prime},q^{\prime\prime})\leq(p^{\prime},q^{\prime})\leq(p^{\prime},q^{\prime\prime})$. We define a function $r^{\prime\prime}$ with domain $\operatorname{dom}(r^{\prime})$ by induction such that for every $\beta$,
@@
The refinement property is clear.
Regarding the mixing property, let $(p,q_{0},r_{0}),(p,q_{1},r_{1})\leq(p,q,r)$. Choose $p_{0},p_{1}\leq p$ such that $p_{0}(1)$ and $p_{1}(1)$ are incompatible and find $q^{\prime}$ such that $(p,q^{\prime})\leq(p,q)$ and $(p_{i},q^{\prime})\leq(p,q_{i})$. It follows that for any inaccessible cardinal $\gamma$, $(p_{0}\upharpoonright\gamma,q^{\prime}\upharpoonright\gamma)$ and $(p_{1}\upharpoonright\gamma,q^{\prime}\upharpoonright\gamma)$ are incompatible. Define a function $r^{\prime}$ on $\operatorname{dom}(r_{0})\cup\operatorname{dom}(r_{1})$ by induction on $\beta$ such that $(p\upharpoonright\beta,q^{\prime}\upharpoonright\beta,r^{\prime}\upharpoonright\beta)\leq(p,q\upharpoonright\beta,r\upharpoonright\beta)$ and $(p_{i}\upharpoonright\beta,q^{\prime}\upharpoonright\beta,r^{\prime}\upharpoonright\beta)\leq(p\upharpoonright\beta,q_{i}\upharpoonright\beta,r_{i}\upharpoonright\beta)$. Assume $\beta\in\operatorname{dom}(r_{0})\cup\operatorname{dom}(r_{1})$ (if $\beta$ is only in one domain, set $r_{i}(\beta)=\emptyset$ for the other $i$). By the inductive hypothesis, $(p_{i}\upharpoonright\beta,q^{\prime}\upharpoonright\beta,r^{\prime}\upharpoonright\beta)$ are incompatible for $i\in 2$, so we can choose a name $r^{\prime}(\beta)$ that is forced by $(p_{0}\upharpoonright\beta,q^{\prime}\upharpoonright\beta,r^{\prime}\upharpoonright\beta)$ to be equal to $r_{0}(\beta)$ and by conditions incompatible with $(p_{0}\upharpoonright\beta,q^{\prime}\upharpoonright\beta,r^{\prime}\upharpoonright\beta)$ to be equal to $r_{1}(\beta)$. Then $(p\upharpoonright\beta+1,q^{\prime}\upharpoonright\beta+1,r^{\prime}\upharpoonright\beta+1)\leq(p\upharpoonright\beta+1,q\upharpoonright\beta+1,r\upharpoonright\beta+1)$ and $(p_{i}\upharpoonright\beta+1,q^{\prime}\upharpoonright\beta+1,r^{\prime}\upharpoonright\beta+1)\leq(p\upharpoonright\beta+1,q_{i}\upharpoonright\beta+1,r_{i}\upharpoonright\beta+1)$. ∎
@@
Let $\nu<\kappa$ be an inaccessible cardinal such that $l\upharpoonright\nu\colon\nu\to V_{\nu}$. Let $\xi:=\nu+1$ and $G$ be $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-generic. In $V[G]$, define the partial order $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi,\beta)$ by induction on $\beta\geq\xi$, letting $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi,\xi):=\{\emptyset\}$ (this defines a name $\mathbb{M}_{3}^{l}(\Gamma,\tau,\mu,\kappa\smallsetminus\xi,\beta)$) Assume that for all $\gamma\in[\xi,\beta)$, $\mathbb{M}_{3}^{l}(\Gamma,\tau,\mu,\kappa\smallsetminus\xi,\gamma)$ has been defined and there is a dense embedding from $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma)$ into $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)*\mathbb{M}_{3}^{l}(\Gamma,\tau,\mu,\kappa\smallsetminus\xi,\gamma)$. Then we let $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi,\beta)$ consist of tripels $(p,q,r)$ such that
@@
    $q$ is a $\mu-$sized partial function on $\beta\smallsetminus\xi$ such that for all $\gamma\in\operatorname{dom}(q)$, $\gamma$ is the double successor of an inaccessible cardinal $\delta$ and $q(\gamma)$ is an $\operatorname{Add}(\tau,\gamma\smallsetminus\delta)$-name for an element in $\dot{\operatorname{Coll}}(\check{\mu},\check{\delta})$.
@@
    $r$ is a $\mu-$sized partial function on $\beta\smallsetminus\xi$ such that for all $\gamma\in\operatorname{dom}(r)$, $\gamma$ is inaccessible, $l(\gamma)$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma)$-name for a $\gamma-$directed closed partial order and $r(\gamma)$ is an $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi,\gamma)$-name for an element in $l(\gamma)^{G}$ (reimagining $l(\gamma)$ as an $\mathbb{M}(\tau,\mu,\xi)*\mathbb{M}_{3}(\Gamma,\tau,\mu,\kappa\smallsetminus\xi,\gamma)$-name).
@@
We do the proof by induction on $\beta\in[\xi,\kappa]$ with the base case being clear. Assume that such an embedding exists for all $\gamma\in[\xi,\beta)$. So in particular, the $\mathbb{M}_{3}^{l}(\Gamma,\tau,\mu,\kappa\smallsetminus\xi,\beta)$ is defined. We define the following embedding:
@@
where $\overline{q}$ and $\overline{r}$ are defined as follows:
$\overline{q}$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-name for a function such that $\Vdash``\operatorname{dom}(\overline{q})=\check{\operatorname{dom}(q)\smallsetminus\xi}^{\prime\prime}$ and for any $\gamma\in\operatorname{dom}(q)\smallsetminus\xi$, $\overline{q}(\check{\gamma})$ is an $\operatorname{Add}(\tau,\xi)$-name (and thus an $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-name) for an $\operatorname{Add}(\tau,\gamma\smallsetminus\xi)$-name corresponding to the $\operatorname{Add}(\tau,\gamma)$-name $q(\xi)$.
$\overline{r}$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-name for a function such that $\Vdash\operatorname{dom}(\overline{r})=\check{\operatorname{dom}(r)\smallsetminus\xi}$ and for $\gamma\in\operatorname{dom}(r)\smallsetminus\xi$, $\overline{r}(\check{\gamma})$ is an $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-name for an $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma\smallsetminus\xi)$-name corresponding to the $\mathbb{M}_{3}^{l}(\tau,\mu,\gamma)$-name $r(\gamma)$ (using the inductive hypothesis).
It follows as in Lemma 5.6 that $\iota$ is a dense embedding. ∎
@@
    $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$ is $\tau-$directed closed,
@@
    The term ordering on $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$ is $\mu-$closed,
@@
Let $\xi=\nu+1$ for an inaccessible cardinal $\nu$. Let $G$ be $\mathbb{M}_{3}^{l}(\tau,\mu,\xi)$-generic. In $V[G]$, let $\dot{\mathbb{L}}$ be an $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)$-name for a $\kappa-$directed closed partial order. Then $\mathbb{M}_{3}^{l}(G,\tau,\mu,\kappa\smallsetminus\xi)*\dot{\mathbb{L}}$ has the $\tau^{+}-$approximation property.